OpenStudy (anonymous):

If f(x) = x+3/3 , then what is f^–1(x) ?

myininaya (myininaya):

is that (x+3)/3 or x+3/3 ?

myininaya (myininaya):

I bet you mean (x+3)/3

OpenStudy (anonymous):

It's supposed to be a fraction, x+3 over 3

myininaya (myininaya):

So this function you have is linear and therefore 1 to 1. Which means it has an inverse. To find the inverse first I will ask you to solve the following for x \[y=\frac{x+3}{3}\]

myininaya (myininaya):

You want to get x by itself.

OpenStudy (anonymous):

Would I divide the 3 on top with the 3 on bottom to get the x by itself?

myininaya (myininaya):

No... You want to undo the division by 3 So you need to multiply both sides by what?

OpenStudy (anonymous):

1/3?

myininaya (myininaya):

No. If we multiply by 3 then wouldn't that 3 cancel with the 3 on bottom?

myininaya (myininaya):

But if you multiply 3 on one side, you must do it to the other

OpenStudy (anonymous):

wouldn't that make it x+9/9 ?

myininaya (myininaya):

Look it is being divided by 3 \[y=\frac{x+3}{3}\] (see the 3 on bottom that means it is being divided by that 3) To undo division by 3 we multiply by 3 like so: \[3 \cdot y=\frac{x+3}{3} \cdot 3 \]

myininaya (myininaya):

Now what happens if you have 3/3?

OpenStudy (anonymous):

it's a 1?

myininaya (myininaya):

Right! 3/3=1 since 1*3=3 So we have \[3y=x+3 \] Now you can finish solving for x?

OpenStudy (anonymous):

x=3y-3?

myininaya (myininaya):

right! now replace x with f^(-1)(x) and y with x

OpenStudy (anonymous):

So the answer ends up being f^(-1)(x)=3x-3?

myininaya (myininaya):

yes

OpenStudy (anonymous):

yayyy thank you for the help :)

myininaya (myininaya):

np :)

myininaya (myininaya):

thanks for your hard work! :)